OpenStudy (anonymous):

if log(1234) = 3.0913 , log(769874)=5.88642) , find the value of \[\sqrt[8]{0.000000001234}\]

OpenStudy (anonymous):

@satellite73 @FoolForMath

hero (hero):

Hint: \[\log_b(y) = x \equiv b^x = y\]

OpenStudy (anonymous):

@Hero No idea!

OpenStudy (anonymous):

is that first one log base ten?

OpenStudy (anonymous):

yes all are with base 10

hero (hero):

b = 10

OpenStudy (anonymous):

then \[\log(\sqrt[8]{.000000001234})=\frac{1}{8}\log(1234\times 10^{-12})\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

check that i counted the decimal places correctly

OpenStudy (anonymous):

u r correct

hero (hero):

Is FFM still typing?

hero (hero):

aka Across

OpenStudy (anonymous):

you get \[\frac{1}{8}\left(\log(1234)+\log(10^{-12})\right)\] \[\frac{1}{8}\left(\log(1234)-12\right)\]

OpenStudy (anonymous):

\[ x =\sqrt[8]{0.000000001234}= (1234 \times 10^{-12})^{\large \frac18} \] \[ \implies \log x = \frac 18 \times (\log 1234 -12 ) \]

OpenStudy (anonymous):

hard for me to count that high, but if @FoolForMath got the same thing, then i guess it must be right

OpenStudy (anonymous):

yup"

OpenStudy (anonymous):

so wat is up next!

OpenStudy (anonymous):

nothing you are done replace \(\log(1234)\) by 3.0913

OpenStudy (anonymous):

i miss across

OpenStudy (anonymous):

Plzz help

OpenStudy (anonymous):

what do you need?

OpenStudy (anonymous):

the answer should be 0.0769874

OpenStudy (anonymous):

after solving urs i got -1.11358

hero (hero):

I got 0.0769874 as well

OpenStudy (anonymous):

then plz show it lol

OpenStudy (anonymous):

i don't know. \[\frac{1}{8}\left(\log(1234)-12\right)=\frac{1}{8}\left(3.0913-12\right)\] then a calculator

hero (hero):

1/8 of a negative number?

OpenStudy (anonymous):

@satellite73, we don't need a calculator, \[\log x = \frac 18 (3.0913-12) = -1.11359 =-2 + 0.88642\] \[\implies x= 10^{-2} \times \text{ Antilog}( 0.88642)=0.0769874 \]

hero (hero):

You only came up with that after the fact

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you are telling me i want to compute \[\log x = \frac 18 (3.0913-12) = -1.11359 =-2 + 0.88642\] with pencil and paper?

OpenStudy (anonymous):

Yes, that's what we are supposed to do in exam :)

OpenStudy (anonymous):

This is a typical entrance problem.

OpenStudy (anonymous):

that is why i do not take exams. but lets say for a minute that you actually managed to do that explain how i am going to compute \[ x= 10^{-2} \times \text{ Antilog}( 0.88642)=0.0769874\]

hero (hero):

lol

hero (hero):

yeah, without a calculator @FFM

OpenStudy (anonymous):

thzz

OpenStudy (anonymous):

By reading the question once again.